OpenStudy (anonymous):

Help me please ! 1. lnabc= 2.log3x(x^3+1)= 3. log6 sqrt(a) (a-1) = 4.2ln7+ln(x+5)-ln(x+3)

OpenStudy (anonymous):

is there a question for these?

OpenStudy (anonymous):

Use the properties of logarithms to expand the following logarithm.

OpenStudy (anonymous):

oooh ok is it \(\ln(abc)\)?

OpenStudy (anonymous):

if so, you can expand as \[\ln(abc)=\ln(a)+\ln(b)+\ln(c)\] others are similar

OpenStudy (anonymous):

k. thanks